Please help. Please provide answer and formula.

Styles Paragraph LU 11-2 Dase of Agil 10 540.000 Jene 10 CHAPTER ORGANIZER AND STUDY GUIDE WITH CHECK FIGURES FOR EXTRA PRACTICE QUIZZES Key polet, procedore, formals Examplels) to ilustrate situation Topic Simple discount note p280 count- MV x iscount x Time $6.000 09瓃.seo leterest based on amount paid back and and pays back 56,000 at maturity ahe not what ceived 60 days A Treasury bill is a good esample of smole oncount note ae $10,000 note, discount rate 12% to 60 days Effective rae p2 2 -$200 What borrower ces Face value Dscount -1224% 9 800 Amount brrower receved Discounting an nerest-eaing noe, p 282 dated Novembr and diacote Deretter at 8% - 51 000 x 0$10 2: 30 4 $1.030 $1.003.27
Answer #1

Ordinary Interest Note: Length of Interest Bank Discount Date of Date of Note Face Value note (Days) rate Rate discount 10-Apr $40,000 170 5% 2%) 10-Jun Interest Face Value Interest rate Time $40,000*5%*(I70/360) 10 $944.44 =D4*F4(E4/360) Maturity Value Face Value Interest $40,000+$944.44 40,944.44 D4+D10 12 14 15 16 17 Discount period is the period between the date of discount and the due date 19 Date of discount 10-Jun Date of note 10-Apr Length of note 170 days Due date -Date of Note +Length of note 2б 27 27-Sep D22+D23 Discount period -Due Date - Date of Discount 29 109 D26-D20 Hence Discounting period is 31 109 Days 32 Maturity Value*Bank Discount Rate*(Discount Period/ 360) $40,944.44*296*(109/360) Bank Discount 35 36 37 38 39 $247.94 D14*G4*(D31/360) Hence Bank Discount is$247.94 41 42 43 Proceeds Maturity Value - Bank Discount $40,944.44 - $247.94 S40,696.50 47 Hence proceeds is $40,696.50

Formula sheet

A B C D E F G H I
2 Ordinary Interest Note:
3 Date of Note Face Value Length of note (Days) Interest rate Bank Discount Rate Date of discount
4 43565 40000 170 0.05 0.02 43626
5
6 a)
7
8 Interest =Face Value*Interest rate*Time
9 =$40,000*5%*(170/360)
10 =D4*F4*(E4/360) =D4*F4*(E4/360)
11
12 Maturity Value =Face Value + Interest
13 =$40,000+$944.44
14 =D4+D10 =D4+D10
15
16 b)
17
18 Discount period is the period between the date of discount and the due date.
19
20 Date of discount =H4
21
22 Date of note =C4
23 Length of note 170 days
24
25 Due date =Date of Note + Length of note
26 =D22+D23 =D22+D23
27
28 Discount period =Due Date - Date of Discount
29 =D26-D20 =D26-D20
30
31 Hence Discounting period is =D29 Days
32
33 c)
34
35 Bank Discount =Maturity Value*Bank Discount Rate*(Discount Period / 360)
36 =$40,944.44*2%*(109/360)
37 =D14*G4*(D31/360) =D14*G4*(D31/360)
38
39 Hence Bank Discount is =D37
40
41 d)
42
43 Proceeds =Maturity Value - Bank Discount
44 =$40,944.44 - $247.94
45 =D14-D39
46
47 Hence proceeds is =D45
48